### v1

@@ A Further Implementation Details and Algorithms
Here, we provide more implementation details in addition to Section 4.1. In our implementation, we use the following set of learning rates. For $\mathcal{L_{\mathrm{Gaussians}}}$ we use: $lr_{color}=0.005,lr_{scaling}=0.005,lr_{rotation}=0.005$ that regulate speed of updating $c_{i}^{j},s_{i}^{j},r_{i}^{j}$ respectively. For $\mu_{i}^{j}$ we apply Exponential annealing of learning rate from $lr_{xyz}=0.00016$ to $lr^{\prime}_{xyz}=0.0000016$ during all 3000 training steps for the first stage. We update SMPL-X body and shape parameters with $lr_{pose}=0.0002$. For $\mathcal{L_{\mathrm{texture}}}$ we use $lr_{texture}=0.01$ to update texture colors. For $\mathcal{L_{\mathrm{filtering}}}$ we use $lr_{color}=0.005,lr_{opacity}=0.05$ for $c_{i}^{j}$ and $o_{i}^{j}$ respectively.
@@ B Additional Comparisons
Figure 10: Novel poses comparison with 3dgs-Avatar [49]. HAHA generates sharper face details, while 3dgs-Avatar produces sharp avatar edges. Artifacts amount for unseen regions is similar, while HAHA uses four time fewer Gaussians.
Figure 11: Additional results for novel poses for SnapshotPeople dataset. HAHA generates less blurred results with fewer artifacts for high-frequency arias such as faces.
In this section we provide additional qualitative comparisons with 3dgs-Avatar [49] and GART [38]. We qualitatively compare HAHA to 3dgs-Avatar on the SnapshotPeople dataset and provide a comparison for more subjects for GART.
3dgs-Avatar. We provide qualitative comparison (Fig. 10) with 3dgs-Avatar [49] in novel poses for SnapshotPeople [7] dataset (similar to Fig. 5 in the main paper). 3dgs-Avatar produces sharper edges of avatars and almost no artifacts in the unseen regions, but it renders blurred faces and uses four times more Gaussians.
Qualitative results for SnapshotPeople. In the main paper, we provide quantitative results for four people in Table 1 and qualitative results for two people in Figure 5 where we compare with the GART [38]. In Figure 11, we provide an additional qualitative visualization in novel poses for the rest of the people included in Table 1 of the main paper. HAHA generalizes to the novel poses and views producing sharp and consistent details (e.g. belt or face for the top row, edge of sweater for bottom row).
Additional comparisons for X-Humans. We provide additional visualizations for X-Humans [53] dataset in Figure 12. Table 5 shows corresponding quantitative metrics for these subjects. These visualizations demonstrate the same quality as the one in the main paper but provide additional examples of our results.
@@ C Gaussian Density Analysis
Table 5: Quantitative metrics for additional subjects from X-Humans [53] dataset.
00018 (male) 00027 (female)
@@
GART [38] 50,47k 30.98 0.9595 0.0683 47,18k 26.56 0.9449 0.0595
GaussianAvatar [20] 191,58k 29.92 0.9588 0.0744 191,58k 25.69 0.9481 0.0543
HAHA(Ours) 18,57k 31.10 0.9630 0.0579 15,50k 27.26 0.9513 0.0473

### v2

@@ A Further Implementation Details and Algorithms
Here, we provide more implementation details in addition to Section 4.1. In our implementation, we use the following set of learning rates. For $\mathcal{L_{\mathrm{Gaussian}}}$ we use: $lr_{color}=0.005,lr_{scaling}=0.005,lr_{rotation}=0.005$ that regulate speed of updating $c_{i}^{j},s_{i}^{j},r_{i}^{j}$ respectively. For $\mu_{i}^{j}$ we apply Exponential annealing of learning rate from $lr_{xyz}=0.00016$ to $lr^{\prime}_{xyz}=0.0000016$ during all 3000 training steps for the first stage. We update SMPL-X body and shape parameters with $lr_{pose}=0.0002$. For $\mathcal{L_{\mathrm{texture}}}$ we use $lr_{texture}=0.01$ to update texture colors. For $\mathcal{L_{\mathrm{filtering}}}$ we use $lr_{color}=0.005,lr_{opacity}=0.05$ for $c_{i}^{j}$ and $o_{i}^{j}$ respectively.
@@ B Additional Comparisons
Figure 10: Additional visualizations for novel poses for SnapshotPeople dataset. HAHA generates less blurred results with fewer artifacts for high-frequency arias such as faces.
In this section we provide additional qualitative results for SnapshotPeople [7] and X-Humans [53] datasets. We provide a comparison for more subjects from these datasets and additional quantitative metrics for SnapshotPeople.
Qualitative results for SnapshotPeople. Here we provide quantitative (Table 5) and qualitative (Fig. 10) results for more people from SnapshotPeople dataset. In Figure 10, we provide qualitative visualization in novel poses for two more people commonly used for comparison. All previous state-of-the-art methods use these four subjects (Fig. 6, 10) for comparison because they have SMPL fits provided by AnimNerf [11]. HAHA generalizes to the novel poses and views producing sharp and consistent details (e.g. belt or face for the top row, edge of sweater for bottom row).
Qualitative results for X-Humans. We provide additional visualizations for X-Humans [53] dataset in Figure 11. These visualizations demonstrate renders for two more people (00018 and 00027) from Table 1. The first two people are visualized in Figure 5 in the main paper. Presented here qualitative results confirm the conclusion provided in the main paper.
@@ C Gaussian Density Analysis
Table 5: Quantitative metrics for two more people from SnapshotPeople [7] dataset.
male-4-casual female-4-casual
@@
GART [38] 20.37k 33.17 0.9719 0.0376 21.22k 34.42 0.9706 0.0328
HAHA(Ours) 14.64k 27.08 0.9505 0.0432 11.96k 31.15 0.9589 0.0283